OpenStudy (anonymous):

write a rule for linear function in the table. x F(x) -3 -1 0 2 3 5 6 8

Nnesha (nnesha):

this is slope y-intercept form equation y = mx + b where m is slope which is always with x and b is y-intercept

Nnesha (nnesha):

y-intercept is always on y -axis where x =0

Nnesha (nnesha):

you have to find slope slope = rise /run formula to find slope is \[\huge\frac{ y _{2}-y _{1} }{ x _{2}-x _{1}}\] you replace x and y by their values and solve

Nnesha (nnesha):

let me know if you have a question or if you didn't understand ?